OpenStudy (anonymous):

Use Cramer's Rule to solve the linear system: x-y=7 3x+2y=6

OpenStudy (anonymous):

happy feet!!!!!!! can you help me out with a problem

OpenStudy (loser66):

where are you so far?

OpenStudy (anonymous):

I don't know where to start

OpenStudy (anonymous):

I knwow cramer's rule

OpenStudy (anonymous):

but, I haven't used cramer's rule to solve linear expressions yet

OpenStudy (loser66):

the matrix you can form from that is \[A = \left[\begin{matrix}1&-1\\3&2\end{matrix}\right]\] and b = \(\left[\begin{matrix}7\\6\end{matrix}\right]\) tell me det (A) =?

OpenStudy (anonymous):

ok, this is what you do

OpenStudy (anonymous):

1(2)-(-1)(3)=2+3=5

OpenStudy (loser66):

yup

OpenStudy (anonymous):

so det(a) equal 5, right?

OpenStudy (loser66):

now, replace b into the first column of A, you get a new matrix, let say it 's A1, calculate det (A1) , what do you have?

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

cause 7(2)-(-1)(6)=20

OpenStudy (loser66):

yup, then det(A1)/det(A) = ?

OpenStudy (anonymous):

4 :)

OpenStudy (loser66):

hihihi, that's x. So, you get x =4 now , do the same, but this time, replace b into the second column of A, let say it's A2, take its det, and det(A2)/detA you get y =

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

:)?

OpenStudy (loser66):

you see, your problem is